OpenStudy (anonymous):

5. A family is having a pool built in their backyard. If their yard is rectangular and measures 7x by 6x and the pool is circular with a radius of 2x how much of the yard will be left over after the pool is built? (1 point) 38x2 42x2 – 4x2 42x2 – 2x2 38x2

OpenStudy (anonymous):

\[ A_{ground} = A_{yard}-A_{pool} \\ A_{yard} = (7x)(6x) \\ A_{pool} = \pi (2x)^2 \\ A_{ground} = (7x)(6x)- \pi (2x)^2 \]Simplify.

OpenStudy (anonymous):

42x2-4x2?